5. Suppose the utility function is given by U(zı,T2) = 14 min{2x, 3y). Calculate the optimal consumption bundle if income is m, and prices are pi, and p2
Answer #1

At optimality 2X=3Y

ans we also know budget line P1X+P2Y=M

putting X=3/2Y into budget line

P1(1.5Y)+P2Y=M

Y=M/(1.5P1+P2) and X=1.5M/1.5P1+P2